# A Systematic Search of Distant Superclusters with the Subaru Hyper Suprime-Cam Survey

### v1

@@ 5.1 Cluster richness and BCG stellar mass
We have further examined the distributions in finer redshift bins ($\Delta z=0.1$), and found that only in two out of five bins, the $p$-value of the All is less than 0.01. However, we note that the uncertainty in stellar mass measurements would generally be (much) larger than 0.1 dex; this, coupled with the fact that the differences are only at 0.02 dex level (c.f. the large scatter of BCG mass at fixed richness as shown in Figure 17) imply that the BCG stellar mass has at most a marginal dependence on the supercluster environment.
@@ 5.2 BCG color and SFR distributions
We further investigate the possible dependence of rest-frame color and SFR of BCGs on different global environments. For the rest-frame color, we adopt rest-frame magnitudes derived by the template fitting-code MIZUKI (Tanaka, 2015; Tanaka et al., 2018). In addition to the cluster selection criteria mentioned above, we further require the reduced chi-square of the best fit model reduced chisq $<5$ as recommended by Nishizawa et al. (2020) when computing the rest-frame color. The SFR used here is derived by DEmP (described in section 2.3). It is found that both color and SFR of BCGs do not seem to depend on richness. We therefore do not perform any fitting to the distributions; rather, we simply compare the color and SFR distributions of SC-Cls and Iso-Cls using the KS test. Figures 20 and 21 show the distributions. The median values and their uncertainties, as well as the KS test $p$-values of each redshift bin and sample set are presented in Table 5. In general, we find no significant difference in both rest-frame color and SFR between SC-Cls and Iso-Cls suggested by the median and $p$-values.
@@ 6 Discussion and Summary
In this work, we have adopted a physically-motivated definition of superclusters proposed by Chon et al. (2015) that superclusters should eventually collapse in an Universe dominated by dark energy. This definition is preferable as it relies on the hierarchical nature of structure formation with fewer subjective decisions (e.g., the choice of the linking length) in the process of identifying superclusters. Furthermore, the supercluster catalogs built following this definition facilitate a more straightforward comparison among different works. Our supercluster finding method is dedicated to fulfilling this definition with the aid of $N$-body simulations.
@@
Our study here suggests that the evolution of BCGs is not significantly influenced by large-scale structures at the supercluster level below redshift $z=1$. An analogous study was conducted by Luparello et al. (2015), who investigated the dependence of the properties of BGGs on global environments. They found that late-type BGGs tend to be redder, more massive, and have lower SFR in supercluster environments, while early-type BGGs show no dependency on global environments. Given that BCGs of CAMIRA clusters presumably are red, early-type galaxies, our findings are consistent with their results (on BGGs). It will be interesting to investigate how supercluster environments affect other member galaxies. However, this is beyond the scope of the current study.
@@
We acknowledge support from the National Science and Technology Council of Taiwan under grants MOST 111-2112-M-001-043, NSTC 111-2628-M-002-005-MY4, and NSTC 112-2112-M-001-061. This work was supported by JSPS KAKENHI Grant JP22H01260. We are grateful to Shogo Ishikawa for providing mock catalogs for analysis. TCC thanks Sut Ieng Tam and Guan-Ming Su for inspiring discussions, and YHC, YTC and TLC for constant support. YTL thanks IH, LYL and ALL for constant encouragement and inspiration. HYS acknowledges support from NTU Academic Research-Career Development Project under Grant No. NTU-CDP- [^1] L7779. The Hyper Suprime-Cam (HSC) collaboration includes the astronomical communities of Japan and Taiwan, and Princeton University. The HSC instrumentation and software were developed by the National Astronomical Observatory of Japan (NAOJ), the Kavli Institute for the Physics and Mathematics of the Universe (Kavli IPMU), the University of Tokyo, the High Energy Accelerator Research Organization (KEK), the Academia Sinica Institute for Astronomy and Astrophysics in Taiwan (ASIAA), and Princeton University. Funding was contributed by the FIRST program from the Japanese Cabinet Office, the Ministry of Education, Culture, Sports, Science and Technology (MEXT), the Japan Society for the Promotion of Science (JSPS), Japan Science and Technology Agency (JST), the Toray Science Foundation, NAOJ, Kavli IPMU, KEK, ASIAA, and Princeton University. This paper is based on data collected at the Subaru Telescope and retrieved from the HSC data archive system, which is operated by Subaru Telescope and Astronomy Data Center (ADC) at NAOJ. Data analysis was in part carried out with the cooperation of Center for Computational Astrophysics (CfCA) at NAOJ. We are honored and grateful for the opportunity of observing the Universe from Maunakea, which has the cultural, historical and natural significance in Hawaii. This paper makes use of software developed for Vera C. Rubin Observatory. We thank the Rubin Observatory for making their code available as free software at http://pipelines.lsst.io/.

### v2

@@ 5.1 Cluster richness and BCG stellar mass
We have further examined the distributions in finer redshift bins ($\Delta z=0.1$), and found that only in one out of five bins, the $p$-value of the All is less than 0.01. However, we note that the uncertainty in stellar mass measurements would generally be (much) larger than 0.1 dex; this, coupled with the fact that the differences are only at 0.02 dex level (c.f. the large scatter of BCG mass at fixed richness as shown in Figure 17) imply that the BCG stellar mass has at most a marginal dependence on the supercluster environment.
@@ 5.2 BCG color and SFR distributions
We further investigate the possible dependence of rest-frame color and SFR of BCGs on different global environments. For the rest-frame color, we adopt rest-frame magnitudes derived by the template fitting-code MIZUKI (Tanaka, 2015; Tanaka et al., 2018). In addition to the cluster selection criteria mentioned above, we further require the reduced chi-square of the best fit model reduced chisq $<5$ as recommended by Nishizawa et al. (2020) when computing the rest-frame color. The SFR used here is derived by DEmP (described in section 2.3). It is found that both color and SFR of BCGs do not seem to depend on richness. We therefore do not perform any fitting to the distributions; rather, we simply compare the color and SFR distributions of SC-Cls and Iso-Cls using the KS test. Figures 20 and 21 show the distributions. The median values and their uncertainties of each redshift bin and sample set are presented in Table 5. In general, we find no significant difference in both rest-frame color and SFR between SC-Cls and Iso-Cls suggested by the median and $p$-values.
@@ 6 Discussion and Summary
In this work, we have adopted a physically-motivated definition of superclusters proposed by Chon et al. (2015) that superclusters should eventually collapse in an Universe dominated by dark energy. This definition is preferable as it relies on the hierarchical nature of structure formation with fewer subjective decisions (e.g., the choice of the linking length) in the process of identifying superclusters. Furthermore, the supercluster catalogs built following this definition facilitate a more straightforward comparison among different works. Our supercluster finding method is dedicated to fulfilling this definition with the aid of $N$-body simulations. In the following, we first discuss our optimized linking length, which reflects the limit of gravitational bound structures (Section 6.1). We then show that the superclusters detected here are equivalent to the supercluster cores mentioned in the literature (Section 6.2). Finally, we interpret our analysis of BCGs in different global environments (Section 6.3).
6.1 On the Limit of Gravitational Bound Structures
@@ 6.3 On the BCGs in Superclusters
Our study here suggests that the evolution of BCGs is not significantly influenced by large-scale structures at the supercluster level below redshift $z=1$. An analogous study was conducted by Luparello et al. (2015), who investigated the dependence of the properties of BGGs on global environments. They found that late-type BGGs tend to be redder, more massive, and have lower SFR in supercluster environments, while early-type BGGs show no dependency on global environments. Given that BCGs of CAMIRA clusters presumably are red, early-type galaxies, our findings are consistent with their results (on BGGs).
Recently, Einasto et al. (2024) divide local galaxy groups and clusters from Tempel et al. (2014) into four classes based on the group luminosity and study the dependence of the quiescent fraction of BGGs and BCGs on the global environment quantified by the luminosity-density field (Liivamägi et al., 2012). They find that once the group class is fixed, the relative ratio between quiescent galaxies, red star-forming galaxies, and blue star-forming galaxies does not vary with the global environment such as voids, supercluster outskirts, and supercluster cores (see their Table 6). Their results also suggest that the properties of BGGs are mainly shaped by the local environment, which is in agreement with our work at higher redshift. It will be interesting to investigate how supercluster environments affect other member galaxies. However, this is beyond the scope of the current study.
@@
We thank an anonymous referee for insightful comments that have improved the clarity of the manuscript. We acknowledge support from the National Science and Technology Council of Taiwan under grants MOST 111-2112-M-001-043, NSTC 111-2628-M-002-005-MY4, NSTC 112-2112-M-001-061, and NSTC 113-2112-M-001-005. This work was supported by JSPS KAKENHI Grant JP22H01260. We are grateful to Shogo Ishikawa for providing mock catalogs for analysis. TCC thanks Sut Ieng Tam and Guan-Ming Su for inspiring discussions, and YHC, YTC and TLC for constant support. YTL thanks IH, LYL and ALL for constant encouragement and inspiration. HYS acknowledges support from NTU Academic Research-Career Development Project under Grant No. NTU-CDP- [^1] L7779. The Hyper Suprime-Cam (HSC) collaboration includes the astronomical communities of Japan and Taiwan, and Princeton University. The HSC instrumentation and software were developed by the National Astronomical Observatory of Japan (NAOJ), the Kavli Institute for the Physics and Mathematics of the Universe (Kavli IPMU), the University of Tokyo, the High Energy Accelerator Research Organization (KEK), the Academia Sinica Institute for Astronomy and Astrophysics in Taiwan (ASIAA), and Princeton University. Funding was contributed by the FIRST program from the Japanese Cabinet Office, the Ministry of Education, Culture, Sports, Science and Technology (MEXT), the Japan Society for the Promotion of Science (JSPS), Japan Science and Technology Agency (JST), the Toray Science Foundation, NAOJ, Kavli IPMU, KEK, ASIAA, and Princeton University. This paper is based on data collected at the Subaru Telescope and retrieved from the HSC data archive system, which is operated by Subaru Telescope and Astronomy Data Center (ADC) at NAOJ. Data analysis was in part carried out with the cooperation of Center for Computational Astrophysics (CfCA) at NAOJ. We are honored and grateful for the opportunity of observing the Universe from Maunakea, which has the cultural, historical and natural significance in Hawaii. This paper makes use of software developed for Vera C. Rubin Observatory. We thank the Rubin Observatory for making their code available as free software at http://pipelines.lsst.io/.
@@ Appendix A Convergence Test
Appendix B Unweighted purity and completeness
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 23: Supercluster total mass $M_{tot}$ distribution in the unperturbed mock catalog (left column) and perturbed mock catalog (right column). The top and bottom rows show distributions at redshift $z=0.9$ and $z=0.6$, respectively. In each panel, the gray histogram shows the $M_{tot}$ distribution of the ground-truth superclusters. The red dashed histogram and blue dotted histogram are $M_{tot}$ distributions obtained by the mass-weighted linking length and the unweighted linking length, respectively.
As the purity and completeness used for the linking length optimization are weighted by the halo mass (equations 2 and 3), one might be concerned about any potential biases resulted from these metrics. Below we show that the mass distribution of our supercluster sample is unbiased.
We first rerun the optimization process using unweighted purity and completeness. The resulting optimized linking length is systematically shorter than the optimized linking length as shown in Figure 7, but the different is $\lessapprox 0.7\,h^{-1}\,$cMpc. We then further apply the unweighted linking length to the light cones described in Section 3.2. Figure 23 shows the total mass distribution of superclusters at different redshifts in the unperturbed and perturbed light cones. In each panel, the gray histogram is the mass distribution of the ground-truth superclusters. The red dashed and blue dotted histograms are mass distributions of supercluster candidates obtained by mass-weighted linking length and unweighted linking length, respectively. Although using the unweighted metric slightly reduces the optimized linking length, the resulting mass distributions are nearly identical to the mass distributions constructed by the mass-weighted linking length. Furthermore, both supercluster candidate mass distributions appear to be consistent with the ground-truth supercluster mass distribution.
To be more quantitative, we perform a KS test comparing the mass distributions of ground-truth superclusters with supercluster candidates obtained with unweighted and mass-weighted linking lengths. The resulting p-values ($p_{uw}$ for the unweighted metric and $p_{w}$ for the mass-weighted metric) are shown in the texts of each panel, and they suggest these three distributions are consistent with each other when the light cones are not perturbed (left column of Figure 23). We therefore conclude that the supercluster candidates we capture are unbiased regardless of the metric adopted for the optimization process. However, introducing photometric redshift uncertainties might marginally reduce the consistency as shown in the right column of Figure 23.
Appendix C Photometric Redshift Confirmation
Here we present more details on our photo-$z$ examination process described in Section 2. We begin by justifying the hyperparameter choice for the DESI examination in Section C.1. Then, using the cluster redshift derived from DESI EDR, we show that the spectroscopic redshift of a BCG can be regarded as a unbiased proxy of the redshift of its host cluster (Section C.2). Lastly, in Section C.3, we validate the hyperparameter selection for the DEmP photo-$z$ examination.
C.1 DESI
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 24: Upper left panel: The confirm fraction of CAMIRA clusters in the DESI SV3 footprint validated by DESI EDR as a function of redshift interval size $\Delta z$ (see Section 2.2 for detail). The curves of different colors represent different minimum DESI galaxies required in our DESI examination process. Upper right panel: similar to the upper left panel, but for the bias between CAMIRA photo-$z$ $z_{CAM}$ and cluster redshift derived from DESI EDR $z_{cl,DESI}$. Lower left panel: similar to the upper left panel, but for the scatter between CAMIRA photo-$z$ $z_{CAM}$ and cluster redshift derived from DESI EDR $z_{cl,DESI}$. Lower right panel: a comparison between spectroscopic redshift of BCGs $z_{BCG}$ and $z_{cl,DESI}$. The bias and scatter are calculated by equations 5 and 6, respectively. Outlier rate (OLR) is the fraction of redshift $4\sigma-$clipped when calculating the scatter as defined in Oguri et al. (2018).
For Section 2.2, we need to determine the appropriate $\Delta z$ and $N_{min}$ values by conducting a parameter search. We first collect $1035$ high-$z$ ($z=0.5-1.0$) CAMIRA clusters located in the one-percent survey footprint of DESI (DESI Collaboration et al., 2023, hereafter SV3), which has the highest completeness, as our test sample. We then repeat the procedure described in Section 2.2 for different $\Delta z$ and $N_{min}=3,4,5$. The upper left panel of Figure 24 shows the confirmed cluster fraction as a function of $\Delta z$ and $N_{min}$. The upper right and lower left panels show the bias and scatter as a function of $\Delta z$ and $N_{min}$. The bias and scatter are calculated using the equations 5 and 6 where $z_{BCG}$ is replaced by $z_{cl,DESI}$.
We see that both the confirmed fraction and the scatter increase initially with $\Delta z$ but become saturated when $\Delta z\geq 0.02$. The saturation suggests that statistically no new clusters could be confirmed, even if we search for structures within a larger initial volume. The saturated scatter is around $0.008$, which is close to the estimate of Oguri et al. (2018). Therefore, in this work, we set $\Delta z$ to $0.02$.
As for $N_{min}$, the scatter seems to be independent of $N_{min}$ as shown in lower left panel of Figure 24. We estimate the background value by randomly placing $5000$ cylinders with a radius $1h^{-1}\,$pMpc and line-of-sight distance of $\Delta v=\pm 3000\,$km/s in the SV3 footprint and find that about $1.3$ galaxies are enclosed within such a cylinder on average. Therefore, we set $N_{min}=4$, which is approximately three times the background value.
C.2 BCG
BCGs are of a special galaxy population residing in galaxy clusters. Although a significant number of BCGs are found in motion relative to their host clusters, the typical peculiar velocity is around $100$ to $200$ km s-1 (Coziol et al., 2009; De Propris et al., 2021), which is much smaller than the cluster photo-$z$ scatter. In Section 2.1, we use the spectroscopic redshift of BCG $z_{BCG}$ as the cluster redshift if it is available and if $|z_{BCG}-z_{CAM}|<0.01(1+z_{CAM})$ where $z_{CAM}$ is the cluster photo-$z$ provided in the CAMIRA catalog.
Here, we investigate whether the redshift of BCGs $z_{BCG}$ is representative of cluster redshift by comparing the available BCG spectroscopic redshift $z_{BCG}$ with clusters validated by DESI EDR (see Section 2.2). The lower right panel of Figure 24 shows the redshift comparison for clusters with available $z_{BCG}$ and are confirmed by our DESI photo-$z$ examination process described above (i.e., $z_{cl,DESI}$ is available). The bias and scatter are calculated by equations 5 and 6 with $z_{CAM}$ replaced by $z_{BCG}$ and $z_{BCG}$ replaced by $z_{cl,DESI}$. The small bias and scatter are $\approx-0.0001(1+z_{cl,DESI})$ and $\approx 0.0015(1+z_{cl,DESI})$, suggesting using $z_{BCG}$ to represent cluster’s redshift is robust.
C.3 DEmP
Denominator Numerator
Purity Richness $N_{mem}\geq N_{min}$ In SV3 footprint Validated by DEmP examination All conditions in the purity denominator Validated by DESI examination
Completeness Richness $N_{mem}\geq N_{min}$ Validated by DESI examination All conditions in the completeness denominator Validated by DEmP examination
Table 7: Definition of purity and completeness in Figure 25
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 25: The confirmed count, purity, and completeness as a function of hyperparameters $(\Delta z,w_{thres},N_{min})$ for DEmP examination described in Section 2.3. Top row: the confirmed count of CAMIRA clusters as a function of the minimum member $N_{min}$. Curves of different colors represent different membership probability thresholds $w_{thres}$ adopted. Middle row: same as the top row, but for the purity defined in Table 7. Bottom row: same as the top row, but for the completeness defined in Table 7. The size of redshift interval $\Delta z=0.01$ and $\Delta z=0.02$ are used for the left and right column, respectively.
Here we conduct a parameter search in order to find the most suitable hyperparameters $(\Delta z,w_{thres},N_{min})$ for the DEmP examination described in Section 2.3. Figure 25 shows the confirmed cluster count, purity, and completeness as a function of $(\Delta z,w_{thres},N_{min})$. The confirmed count is the number of clusters in our whole sample validated by the DEmP examination for different hyperparameter choices. As for the purity and completeness, in table 7, we list the conditions that clusters have to satisfy simultaneously when considering the clusters for denominator and numerator of purity and completeness. In this way, the purity quantifies the fraction of clusters confirmed by the DESI examination among the clusters confirmed by the DEmP examination. The completeness quantifies the fraction of clusters confirmed by the DEmP examination among the clusters confirmed by the DESI examination.
In Figure 25, we first find that when we adopt $\Delta z=0.02$ instead of $\Delta z=0.01$, the purity is slightly reduced, but the completeness is increased dramatically. Then, considering that both purity and completeness are not a strong function of $N_{min}$ (the variation is $\lesssim 10\%$), we adopt $N_{min}=10$ since a higher $N_{min}$ would strongly reduce the confirmed count. Lastly, as we prioritize purity over completeness, we settle the membership probability threshold to $w_{thres}=0.5$, i.e., we only use reliable member galaxies for this analysis. In summary, we set our hyperparameter choice to $(\Delta z,w_{thres},N_{min})=(0.02,0.5,10)$, leading to purity and completeness of $\approx 82\%$ and $\approx 71\%$, respectively.
Appendix D Fragmentation
It is possible that a high-multiplicity supercluster is split or identified as several low-multiplicity superclusters. We investigate this “fragmentation” issue here. We begin our analysis by comparing the ground-truth supercluster catalog with the FoF groups (supercluster candidates) catalog in the simulation boxes. The upper panel of Figure 26 shows the mass recovery ratio as a function of the multiplicity $N_{\rm SC}$ of ground-truth supercluster at redshift $z=1.0$ (left panel) and $z=0.5$ (right panel). For each ground-truth supercluster (represented by a gray point in the Figure), we calculate the mass recovery ratio by picking up their most massive FoF group counterpart, which is a subset of the ground-truth supercluster cluster members. As suggested by the median shown as the red curve, for most ground-truth supercluster of $N_{\rm SC}=2-3$, which dominate the supercluster population (Figure 13), our method can recover their cluster members without splitting them. However, for ground-truth superclusters of $N_{\rm SC}>3$, our method generally recovers $\approx 80\%$ of their mass, reflecting the fragmentation problem. The text in the lower right corner shows the fraction $f_{missed}$ of ground-truth superclusters not detected at all by our method, i.e., the mass recovery ratio equals zero. The value $f_{missed}\approx 20\%$ is consistent with our optimized completeness shown in Figure 8.
The lower panel of Figure 26 shows 2D histograms between the detected multiplicity and the multiplicity of the ground-truth superclusters. Normalization is set such that the sum of value in each column is unity. For example, at redshift $z=1.0$, there is one ground-truth supercluster with $N_{\rm SC}=9$ completely captured. As another example, at redshift $z=0.5$, there is one ground-truth supercluster with $N_{\rm SC}=15$ (the right most column) identified as a $N_{\rm SC}=7$ FoF group, one $N_{\rm SC}=4$ FoF group, and two $N_{\rm SC}=2$ FoF groups. We note that the FoF groups with multiplicity higher than their ground-truth supercluster counterparts are excluded in the analysis. This is because such FoF groups are regarded as contamination and only account for less than $5\%$ of all FoF groups by number. In summary, from the 2D histograms, we see that while the ground-truth superclusters of $N_{\rm SC}\leq 4$ are generally not split, the fragmentation issue becomes worse for ground-truth supercluster of higher multiplicity, as can be expected. Although the high-multiplicity superclusters are extremely rare, this analysis illustrates the limitation of using distance solely to connect structures that will collapse in the future.
In Figure 26, the effect of photo-$z$ uncertainties is not accounted for, but it is expected to worsen the issue. This also explains the absence of high-multiplicity superclusters in Figure 13. This again reveals the importance of using spectroscopic data to precisely locate large-scale structures at higher redshift.
Refer to captionRefer to caption
Figure 26: Left panel: The upper subplot shows the mass recovery ratio as a function of the multiplicity $N_{\rm SC}$ of the ground-truth superclusters. Each gray point represents a ground-truth supercluster in our six simulation boxes. The red curve shows the median as a function of $N_{\rm SC}$. The fraction of ground-truth superclusters that are not detected $f_{missed}$ is shown in the lower right corner. The lower subplot shows a 2D histogram between the detected multiplicity and the ground-truth multiplicity. Normalization is set such that each column sums up to unity. Right panel: same as left panel, but at redshift $z=0.50$.